OpenStudy (anonymous):

Solve the system of equations. 3x - 2y = 9 x + y = 8 (6, 2) (3, 0) (1, 6) (5, 3)

OpenStudy (anonymous):

Fill in the choices 3*6=18-4=14 So no 9-0= 0 yes and 3+0=8 no 3-12=9 no 15-6=9 yes and 5+3=8 yes So the answer is d. (5,3)

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

You are welcome, thank you for the medal!